OpenStudy (anonymous):

ason and Ron together have 97 baseball cards. Jason has four more than twice as many cards as Ron. How many baseball cards does each person have? A. Jason has 31 cards and Ron has 66 cards B. Ron has 31 cards and Jason has 66 cards C. Jason has 46 cards and Ron has 51 cards D. Ron has 46 cards and Jason has 51 cards

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

hey, you didn't answer my question on the last question :-)

OpenStudy (whpalmer4):

what was your answer, and how much did 1 ticket cost?

OpenStudy (anonymous):

My answer was D

OpenStudy (whpalmer4):

good. Jason and Ron together have 97 baseball cards. All of the answer choices have 97 cards, so no shortcut there. Jason has 4 more than twice as many cards as Ron. Which choices show Jason having about twice as many cards as Ron?

OpenStudy (anonymous):

C & D?

OpenStudy (whpalmer4):

Well, no. Let's put on the thinking cap. If Jason has 4 more than twice as many as Ron, Jason's number needs to be bigger than Ron's, right? That would seem to eliminate C as a possible correct answer.

OpenStudy (anonymous):

Okay

OpenStudy (whpalmer4):

And as for D, is 51 about twice as many as 46?

OpenStudy (anonymous):

...LoL. no

OpenStudy (whpalmer4):

Okay. go take a look at the other choices and see if you can figure it out.

OpenStudy (anonymous):

So. Eliminate c and d

OpenStudy (anonymous):

hmmm...Idk. Lol

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (whpalmer4):

:-(

OpenStudy (whpalmer4):

What are the first two choices?

OpenStudy (anonymous):

Is it b?

OpenStudy (whpalmer4):

I don't know what b is. just tell me the number of cards each guy has.

OpenStudy (anonymous):

Ron has 31, and Jason has 66

OpenStudy (whpalmer4):

Okay. Ron has 31. 2*31 = 62. Jason has 4 more than 2x as many as Ron. 2*31 + 4 = 62 + 4 = 66. That wasn't so hard, was it?

OpenStudy (anonymous):

Lol

OpenStudy (whpalmer4):

If we had to solve it without the benefit of multiple choice answers: \[R + J = 97\]\[2R + 4 = J\] Substitute the second equation in the first: \[R + (2R + 4) = 97\]\[3R + 4 = 97\]\[3R = 93\]\[R = 31\]\[J = 2R+4 = 2(31)+4 = 66\]

OpenStudy (whpalmer4):

Gotta run, catch you later...